2) You go to the shoe store and are looking for new running shoes. At the store you find out that they have 32 different pairs of shoes that would fit you. If 12 of those are Nike’s, and 10 are Reebok, what is the probability if you choose randomly that you would get a shoe that is made by Nike or Reebok? Please set up (i.e P(Nike)) and show all steps of the question, not just the answer.

3) A survey was conducted of people ages 18-22 living in Santa Barbara. From this survey it was determined that 68% of them are students, while 31% are working. It is also determined that 15% work and are students at the same time. What is the probability a randomly chosen person between the ages of 18-22 living in Santa Barbara is a student or works? Please set up(i.e. P(Student)) and show all steps of the question, not just the answer.
Answer #1

2) Total pairs of shoes that would fit you= 32

Number of Nike's shoes= 12

P( Nike) = 12/32

Number of Reebok's shoes = 10

P( Reebok) =10/32

P( Nike or Reebok) = P( Nike ) + P( Reebok) ; as Nike and Reebok are two mutually exclusive events that means = (12/32) + (10/32)   that a pair of shoes cant be of Nike and Reebok at the same time  

= 22/32

= 0.6875

3) P( Student) =0.68

P( working) = 0.31

P( Student & working) = 0.15

P( Student OR working) = P( Student) + P( working) - P( Student & working)

= 0.68 + 0.31 - 0.15

= 0.84